The bones of a saber-toothed tiger are found to have an activity per gram of carbon that is 30.0% of what would be found in a similar live animal. How old are these bones?
Answer #1

please give an upvote if you like
Answer #2

SOLUTION :


Radioactive carbon 14C has half-life of 5730 years.


So,


A / A0 = e^(- lamda * t)

=> 1/2 = e^(- lamda * 5730)

Taking ln :

=> ln(1/2) = - lamda * 5730

=> lamda = - ln(1/2) / 5730 = 1.20968 * 10^(-4) 


So,


A(now) / A0 = e^(- 1.20968*10^(-4)  * t)

=> 0.30 =  e^(- 1.20968*10^(-4)  * t)

Takin ln :

=> ln(0.30) = - 1.20968*10^9-4) * t

=> t = ln(0.30) / (- 1.20968*10^(-4)

=> t =  9952.82 years


Hence, the bones are 9952.82 years old (ANSWER)

answered by: Tulsiram Garg